# Yang-Mills extension of the Loop Quantum Gravity-corrected Maxwell equations

### v2

@@ I Introduction
The consistent and precise description of quantum gravity is a big challenge and an important open problem in the theory of fundamental interactions. This quest is going on for more than 50 years, by adopting different scenarios. We point out that some of those try to achieve their goal by also seeking a grand-unified theory. In this framework, one of the most promising theories is Loop Quantum Gravity (LQG) Ashtekar:2004eh ; Ashtekar:2011ni ; Rovelli:2014ssa ; Rovelli:1997yv , whose main purpose is to understand fundamental physics at the Planck scale $l_{P}=10^{-35}$m, through the quantization of gravity and not via a grand-unification approach. The LQG formulation predicts that spacetime must be discrete, granular, at the Planck scale. More precisely, the granularity of spacetime is implemented by means of a mathematically structure referred to as spin foam Rovelli:2014ssa ; Rovelli:1997yv ; Thiemann:2001gmi ; Gambini:2011zz . Moreover, the fields present a background-independent description that manifests itself through an action that exhibits diffeomorphism invariance and is also non-perturbative. There are many ways to describe LQG; in the present article, we choose to work with the canonical method, which is described by Hamiltonians Thiemann:2001gmi . The formalism of LQG is based on the Ashtekar variables Ashtekar:1986yd , introduced in by 1986 Abhay Ashtekar in his formalism to describe General Relativity by an $SU(2)$ gauge connection.
Loop Quantum Gravity offers a consistent theoretical framework which extends its exploration to include phenomena that contribute to the understanding of what gravity is at a fundamental quantum scale. If we aim to detect possible traces of LQG in various events Amelino-Camelia:1997ieq ; Ellis:1999rz ; Amelino-Camelia:1999hpv , the most notable ones are: noise present in gravitational wave detectors Amelino-Camelia:2001dbf ; Amelino-Camelia:1999vks , neutral kaon systems Ellis:1995xd ; Huet:1994kr and the time-dependent energy of arrival of photons and neutrinos from distant sources Alfaro:2001rb ; Li:2022szn . The electromagnetic sector has an important contribution to this. It is expected that at the Planck scale (local) Lorentz invariance, a basic principle of General Relativity, be broken. However, achieving the energy scales necessary to test both Lorentz-invariance violation (LIV) and Planck-scale effects in terrestrial experiments is not feasible Amelino-Camelia:1997ieq ; Li:2022szn ; Gambini:1998it ; Amelino-Camelia:1996bln ; Melo:2024gxl . An alternative is to focus on the energy of Gamma-Ray Bursts (GRBs) and their enormous energy emitted, which may render possible to investigate the energy-dependent time of arrival of photons/neutrinos from distant sources. In this work, we are committed with the photon sector. The speed of light in a vacuum characterized by a granular space-time may assume the form
@@
where $E^{\gamma}_{LIV}\approx 3.6\times 10^{17}$ GeV and is independent of the photon’s helicity. The high-energy scale of the $E^{\gamma}_{LIV}$-parameter supports the idea that GRBs are the most suitable candidates to investigate this parameter  Shao:2009bv ; Zhang:2014wpb ; Xu:2016zxi ; Xu:2016zsa ; Amelino-Camelia:2016ohi ; Amelino-Camelia:2017zva ; Xu:2018ien ; Liu:2018qrg ; Li:2020uef ; Zhu:2021pml ; Chen:2019avc .
@@
In this paper, we are adopting the natural system of units, $c=\hbar=1$ and the signature of Minkowski metric is (+,,,)(+,-,-,-)( + , - , - , - ). Our work is outlined as follows. In Sec. II, we briefly introduce the electromagnetic sector in LQG; for this, we consider the Hamiltonian of the theory and the four modified Maxwell equations in the context of Loop Quantum Gravity. In Section  III, we show the details how, from the Abelian electromagnetic sector with its Hamiltonian and equations of motion, we can obtain a non-Abelian theory of self-interactions for massless vector (gauge) bosons. Finally, in Sec. IV, we present our final comments and discussions.
@@ II LQG Effects in the Electromagnetic Sector
The Hamiltonian formulation of the effects of LQG in the electromagnetic theory was carried out in the work of Ref.  Alfaro:2001rb . This construction was done by considering the spacetime as a manifold $M$ whose topology is a diffeomorphism for $\Sigma\times\mathbb{R}$, where $\Sigma$ is a Riemmannian-3 manifold. The LQG formalism is based on the Ashtekar variables described for a non-compact gauge group, $SL(2,\mathbb{C})$, where the indices $a,b,c$ stand for the space coordinates and $i,j,k$ are their indices assigned to the generators of $su(2)$. The gauge connection is represented as $(A_{a}^{i}=\Gamma_{a}^{i}-iK_{a}^{i})$, with $\Gamma_{a}^{i}$, torsion-free, is a function of the co-triads $e^{i}_{a}$, which in turn are defined as a function of the three-metric $q_{ab}=e^{i}_{a}e^{i}_{b}$. The term $K_{a}^{i}$ is the extrinsic curvature of the Riemmannian 3-manifold, $\Sigma$, and is related to the co-triads according to the relation $K_{ab}=sgn[\mbox{det}(e_{c}^{j})]K_{a}^{i}e_{b}^{i}$. The other variable part of Ashtekar’s formalism is the canonical pair of the conjugate moment, namely, the electric field of the form $E_{i}^{a}=\frac{1}{2}\epsilon^{abc}\epsilon_{ijk}e_{b}^{j}e_{c}^{k}$. The Hamiltonian of the electromagnetic sector modified by LQG is built up based on these considerations, and it takes the form:
$\begin{split}H_{LQG}=\frac{1}{Q^{2}}\int d^{3}x\Biggl{\{}\left[1+\theta_{7}\left(\frac{l_{P}}{\mathcal{L}}\right)^{2+2\Upsilon}\right]\frac{1}{2}(\underline{\vec{B}}^{2}+\underline{\vec{E}}^{2})\\ +\theta_{3}l_{P}^{2}(\underline{B}^{a}\nabla^{2}\underline{B}_{a}+\underline{E}^{a}\nabla^{2}\underline{E}_{a})+\theta_{2}l_{P}^{2}\underline{E}^{a}\partial_{a}\partial_{b}\underline{E}^{b}\\ +\theta_{8}l_{P}[\underline{\vec{B}}\cdot(\nabla\times\underline{\vec{B}})+\underline{\vec{E}}\cdot(\nabla\times\underline{\vec{E}})]\\ +\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(\underline{\vec{B}}^{2})^{2}+...\Biggl{\}},\end{split}$ (2)
where $Q^{2}$ is the coupling constant of electromagnetism, $l_{p}\approx 1.6\times 10^{-35}m$ is the Planck length. The characteristic length, $\mathcal{L}$, satisfies the condition $l_{P}\ll\mathcal{L}\leq\lambda$. The parameter $\lambda$ is the de Broglie wavelength, and the characteristic length, $\mathcal{L}$, has a maximum value at momentum $k$ whenever $\mathcal{L}=k^{-1}$. Another parameter appearing in the equation is $\Upsilon$, which is given by the order of the contribution of the gravitational connection to the expected value and it may be determined by the phenomenological analysis of a selected event. That is, it may depend on the helicity of the particle under consideration Ellis:1999uh ; Ellis:1999yd ; Ellis:1999sd . The $\theta_{i}$’s are non-dimensional parameters of order one or are extremely close to zero Alfaro:2001rb ; Alfaro:2002ya ; finally, $a,b$ are spatial tensor indices. From now on, we replace the space indices by adopting a vector notation, and we eliminate the underline that refers to the canonical pairs for the electromagnetic sector in all quantities. From the Eq. (2), we get the field equations as follows below:
@@
$\begin{split}A_{\gamma}(\nabla\times\vec{B})-\frac{\partial\vec{E}}{\partial t}+2l_{P}^{2}\theta_{3}\nabla^{2}(\nabla\times\vec{B})-2\theta_{8}l_{P}\nabla^{2}\vec{B}\\ +4\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}\nabla\times(\vec{B}^{2}\vec{B})=0,\end{split}$ (4)
@@
$\begin{split}A_{\gamma}(\nabla\times\vec{E})+\frac{\partial\vec{B}}{\partial t}+2l_{P}^{2}\theta_{3}\nabla^{2}(\nabla\times\vec{E})-\\ 2\theta_{8}l_{P}\nabla^{2}\vec{E}=0,\end{split}$ (6)
@@
The Eq. (2) is an infinite expression. In the literature, it is common to truncate it at the first order in the non-linear term in the magnetic field. When the field equations are calculated, this non-linear term contributes to Eq. (4). The main quantities of the electromagnetic theory have already been calculated and analyzed previously, as it can be seen in Alfaro:2001rb ; Li:2020uef . In the present work, such quantities are not of primary interest; the modified Maxwell equations in the context of LQG as well as its corresponding Hamiltonian are the key equations for the development of the present contribution.
@@ III The Noether Procedure as our path towards the non-Abelian scenario
To formulate a non-Abelian theory for massless spin-1 particles, one may follow different approaches. In this paper, we pick out the so-called Noether method Deser:1969wk ; VanNieuwenhuizen:1981ae , which essentially consists in generating self-interactions among vector fields through the variation of the original free Lagrangian and proposing self-interactions by coupling these vector fields with the Noether current of their initially global. To realize this program, we start off from an N-plet of vector fields (at the end of the procedure, they will become our Yang-Mills gauge bosons) in an arbitrary representation of the symmetry group as given below:
$A^{m},m=1,2,...,N\ \ \ \ \ \ \ \phi^{n},n=1,2,...,N$ (8)
in N-dimensions, we can represent these fields in a Lie group that transforms as
@@
We are working in the general case of an $SU(N)$ gauge group. So, $w_{h}$ stands for the parameters of the $SU(N)$ transformations, whereas the 3-index symbol $f_{hij}$ represents the structure constants of $SU(N)$. To initiate the Noether procedure, we start off with the free Lagrangian of the theory. To get this Lagrangian, we perform a Legendre transformation by taking Eq. (2) and writing
$\begin{split}L_{LQG}=\frac{1}{Q^{2}}\int d^{3}x\Biggl{\{}\left[1+\theta_{7}\left(\frac{l_{P}}{\mathcal{L}}\right)^{2+2\Upsilon}\right]\frac{1}{2}(\vec{E}^{2}-\vec{B}^{2})\\ +\theta_{3}l_{P}^{2}(\vec{B}\nabla^{2}\vec{B}+\vec{E}\nabla^{2}\vec{E})+\theta_{2}l_{P}^{2}\vec{E}\cdot[\nabla\cdot(\nabla\vec{E})]\\ +\theta_{8}l_{P}[\vec{B}\cdot(\nabla\times\vec{B})+\vec{E}\cdot(\nabla\times\vec{E})]\\ +\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(\vec{B}^{2})^{2}\Biggl{\}},\end{split}$ (15)
@@
$\begin{split}\vec{j}_{h}=-f_{hij}\{A_{\gamma}[(\vec{E}_{i}\cdot\phi_{j})-(\vec{B}_{i}\times\vec{A}_{j})]-\theta_{3}l_{P}^{2}[(\vec{B}_{i}\times\\ \nabla^{2}\vec{A}_{j})-(\nabla^{2}\vec{B}_{i}\times\vec{A}_{j})-(\nabla^{2}\vec{E}_{i}\cdot\phi_{j})-(\vec{E}_{i}\cdot\nabla^{2}\phi_{j})]\\ -\theta_{8}l_{P}[2(\nabla^{2}\vec{A}_{i}\times\vec{A}_{j})-(\vec{E}_{i}\times\vec{\dot{A}}_{j})-(\nabla\times\vec{E}_{i})\phi_{j}]\\ +4\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(B^{2}\cdot\vec{B}_{i}\times\vec{A}_{j})\},\end{split}$ (18)
@@
$\begin{split}j_{h}^{0}=-f_{hij}\{A_{\gamma}(\vec{E}_{i}\cdot\vec{A}_{j})+\theta_{3}l_{P}^{2}[(\nabla^{2}\vec{E}_{i}\cdot\vec{A}_{j})+\\ (\vec{E}_{i}\cdot\nabla^{2}\vec{A}_{j})]-2\theta_{8}l_{P}(\nabla\times\vec{E}_{i})\cdot\vec{A}_{j}\}.\end{split}$ (19)
@@
The process will be the same, but with an additional step. Before identifying the new conserved currents, we must consult the new field equations derived from the Lagrangian of the previous step. This is so because the new identified currents must be on-shell. This is an essential point. From the Lagrangian, we can derive new Gauss’ electricity and Ampère-Maxwell equations. This allows us to find the currents to be coupled to the fields to once again insert in the Lagrangian. Each current is coupled to a field and depends on a coupling parameter $l^{\prime}$, which represents the coupling constant for this new term
@@
The Eq. (24) is the last Lagrangian expression obtained, and it is no longer necessary to re-calculate other currents. The reason being that the expression of the new Lagrangian no longer depends on derivatives of the fields. Should the same steps be repeated, we would obtain the same expressions for the currents as before. From Eq. (24), we can obtain the field equations, namely, Gauss’ electricity and Ampère-Maxwell. We set $l=\frac{1}{4}g$ and $l^{\prime}=\frac{1}{2}g$ and from these two equations of motion, it is possible to define the equations for the electric and magnetic field equations in terms of the potentials in the Yang-Mills formalism with LQG contributions:
$\begin{split}\vec{E}_{h}=A_{\gamma}[-\nabla\cdot\phi_{h}-\vec{\dot{A}}_{h}+gf_{hij}\phi_{i}\vec{A}_{j}]-\frac{3g}{2}\theta_{3}l_{P}^{2}f_{hij}\cdot\\ [\vec{A}_{i}(\nabla^{2}\phi_{j})+(\nabla^{2}\vec{A}_{i})\cdot\phi_{j}]+\theta_{3}l_{P}^{2}f_{hij}[(\nabla\vec{A}_{i})\cdot(\nabla\phi_{j})]\\ +\frac{g}{2}\theta_{8}l_{P}f_{hij}(\vec{A}_{i}\times\vec{\dot{A}}_{j}),\end{split}$ (25)
@@
$\begin{split}\vec{B}_{h}=A_{\gamma}[(\nabla\times\vec{A}_{h})+\frac{g}{2}f_{hij}(\vec{A}_{i}\times\vec{A}_{j})]+2l_{P}^{2}\theta_{3}\nabla^{2}\vec{B}_{h}\\ -\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\cdot[\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-[\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})]]+\\ \frac{g}{2}\theta_{8}l_{P}f_{hij}[(\vec{\dot{A}}_{i})\cdot\phi_{j}]+2\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}[2(\vec{B}^{2}\vec{B}_{h})\\ +3gf_{hij}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}].\end{split}$ (26)
If we were working with the electromagnetism of the Standard Model of particle physics, all the procedures described so far would be sufficient to obtain all the non-Abelian Maxwell equations. Simply taking the curl of Eq. (25) and the divergence of Eq. (26) would be sufficient. In electrodynamics with LQG effects, this is not possible, because the Faraday-Lenz equation is modified and, by doing this and taking the Abelian limit, we do not obtain again the Eqs. (5), (6) that should be obtained. As it can be consulted in the literature Sorokin:2021tge ; Gaete:2014nda , the non-Maxwellian extensions of electrodynamics does not modify the Faraday-Lenz equation, since together with Gauss’law for magnetism, these equations are obtained from the Bianchi Identity. The electrodynamics presented in this paper modifies Eq. (6), which allows us to conclude that the Bianchi Identities are modified. It is therefore necessary to follow another path to attain these two equations; one possible way consists in getting them from the Hamilton-Jacobi equations. By performing a Legendre transformation on the Lagrangian Eq. (24) to find its corresponding Hamiltonian, we can derive the equations for the electric field, $E$, and the vector potential, $A$:

### v3

@@ I Introduction
The consistent and precise description of quantum gravity is a big challenge and an important open problem in the theory of fundamental interactions. This quest is going on for more than 50 years, by adopting different scenarios. We point out that some of those try to achieve their goal by also seeking a grand-unified theory. Loop Quantum Gravity does not enter this framework because its maajor purpose is to understand fundamental physics at the Planck scale $l_{P}=10^{-35}$m, only the quantization of gravity Ashtekar:2004eh ; Ashtekar:2011ni ; Rovelli:2014ssa ; Rovelli:1997yv . The LQG formulation predicts that spacetime must be discrete, granular, at the Planck scale. More precisely, the granularity of spacetime is implemented by means of a mathematically structure referred to as spin foam Rovelli:2014ssa ; Rovelli:1997yv ; Thiemann:2001gmi ; Gambini:2011zz . Moreover, the fields present a background-independent description that manifests itself through an action that exhibits diffeomorphism invariance and is also non-perturbative. There are many ways to describe LQG; in the present article, we choose to work with the canonical method, which is described by Hamiltonians Thiemann:2001gmi . The formalism of LQG is based on the Ashtekar variables Ashtekar:1986yd , introduced in by 1986 Abhay Ashtekar in his formalism to describe General Relativity by an $SU(2)$ gauge connection.
Loop Quantum Gravity offers a consistent theoretical framework which extends its exploration to include phenomena that contribute to the understanding of what gravity is at a fundamental quantum scale. If we aim to detect possible traces of LQG in various events Amelino-Camelia:1997ieq ; Ellis:1999rz ; Amelino-Camelia:1999hpv , the most notable ones are: noise present in gravitational wave detectors Amelino-Camelia:2001dbf ; Amelino-Camelia:1999vks , neutral kaon systems Ellis:1995xd ; Huet:1994kr and the time-dependent energy of arrival of photons and neutrinos from distant sources Alfaro:2001rb ; Li:2022szn . The electromagnetic sector has an important contribution to this. It is expected that at the Planck scale (local) Lorentz invariance, a basic principle of General Relativity, be broken. However, achieving the energy scales necessary to test both Lorentz-invariance violation (LIV) and Planck-scale effects in terrestrial experiments becomes a significant challenge, although there have been attempts, such as the investigation using of kaon systems Amelino-Camelia:1997ieq ; Li:2022szn ; Gambini:1998it ; Amelino-Camelia:1996bln ; Melo:2024gxl . An alternative is to focus on the energy of Gamma-Ray Bursts (GRBs) and their enormous energy emitted, which may render it possible to investigate the energy-dependent time of arrival of photons/neutrinos from distant sources. In this work, we are committed with the photon sector. The speed of light in a vacuum characterized by a granular space-time may assume the form
@@
where $E^{\gamma}_{LIV}\gtrsim 3.6\times 10^{17}$ GeV and is independent of the photon’s helicity. The high-energy scale of the $E^{\gamma}_{LIV}$-parameter supports the idea that GRBs are the most suitable candidates to investigate this parameter  Shao:2009bv ; Zhang:2014wpb ; Xu:2016zxi ; Xu:2016zsa ; Amelino-Camelia:2016ohi ; Amelino-Camelia:2017zva ; Xu:2018ien ; Liu:2018qrg ; Li:2020uef ; Zhu:2021pml ; Chen:2019avc .
@@
In this paper, we are adopting the natural system of units, $c=\hbar=1$ and the signature of Minkowski metric is (+,,,)(+,-,-,-)( + , - , - , - ). Our work is outlined as follows. In Sec. II, we briefly introduce the electromagnetic sector in LQG; for this, we consider the Hamiltonian of the theory and the four modified Maxwell equations in the context of Loop Quantum Gravity. In Section  III, we show the details how, from the Abelian electromagnetic sector with its Hamiltonian and equations of motion, we can obtain a non-Abelian theory of self-interactions for massless vector (gauge) bosons. To conclude, in Sec. IV, we present our final comments and discussions. The Noether procedure applied to get the Yang-Mills formulation from the classical Maxwellian electromagnetic theory is summarized also given in an Appendix A.
@@ II LQG Effects in the Electromagnetic Sector
The Hamiltonian formulation of the effects of LQG in the electromagnetic theory was carried out in the work of Ref.  Alfaro:2001rb . This construction was done by considering the spacetime as a manifold $M$ whose topology is a diffeomorphism for $\Sigma\times\mathbb{R}$, where $\Sigma$ is a Riemmannian-3 manifold. The LQG formalism is based on the Ashtekar variables described by a non-compact gauge group, $SL(2,\mathbb{C})$, where the indices $a,b,c$ stand for the space coordinates and $i,j,k$ are their indices assigned to the generators of $su(2)$. The gauge connection is represented as $(A_{a}^{i}=\Gamma_{a}^{i}-iK_{a}^{i})$, with $\Gamma_{a}^{i}$, torsion-free, is a function of the co-triads $e^{i}_{a}$, which in turn are defined as a function of the three-metric $q_{ab}=e^{i}_{a}e^{i}_{b}$. The term $K_{a}^{i}$ is the extrinsic curvature of the Riemmannian 3-manifold, $\Sigma$, and is related to the co-triads according to the relation $K_{ab}=sgn[\mbox{det}(e_{c}^{j})]K_{a}^{i}e_{b}^{i}$. The other variable part of Ashtekar’s formalism is the canonical pair of the conjugate moment, namely, the electric field of the form $E_{i}^{a}=\frac{1}{2}\epsilon^{abc}\epsilon_{ijk}e_{b}^{j}e_{c}^{k}$. The Hamiltonian of the electromagnetic sector modified by LQG is built up based on these considerations, and it takes the form:
$H_{LQG}=$ $\frac{1}{Q^{2}}\int d^{3}x\Biggl{\{}\left[1+\theta_{7}\left(\frac{l_{P}}{\mathcal{L}}\right)^{2+2\Upsilon}\right]\frac{1}{2}(\underline{\vec{B}}^{2}+\underline{\vec{E}}^{2})$
[…]
where $Q^{2}$ is the coupling constant of electromagnetism, $l_{p}\approx 1.6\times 10^{-35}m$ is the Planck length. The characteristic length, $\mathcal{L}$, satisfies the condition $l_{P}\ll\mathcal{L}\leq\lambda$. The parameter $\lambda$ is the de Broglie wavelength, and the characteristic length, $\mathcal{L}$, has a maximum value at momentum $k$ whenever $\mathcal{L}=k^{-1}$. Another parameter appearing in the equation is $\Upsilon$, which is given by the order of the contribution of the gravitational connection to the expected value and it may be determined by the phenomenological analysis of a selected event. That is, it may depend on the helicity of the particle under consideration Ellis:1999uh ; Ellis:1999yd ; Ellis:1999sd . The $\theta_{i}$’s are non-dimensional parameters of order one or are extremely close to zero Alfaro:2001rb ; Alfaro:2002ya ; finally, $a,b$ are spatial tensor indices. From now on, we replace the space indices by adopting a vector notation, and we eliminate the underline that refers to the canonical pairs for the electromagnetic sector in all quantities. From the Eq. (II), we get the field equations as follows below:
@@
$A_{\gamma}(\nabla\times\vec{B})-\frac{\partial\vec{E}}{\partial t}+2l_{P}^{2}\theta_{3}\nabla^{2}(\nabla\times\vec{B})-2\theta_{8}l_{P}\nabla^{2}\vec{B}$
$+4\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}\nabla\times(\vec{B}^{2}\vec{B})=0,$ (4)
@@
$A_{\gamma}(\nabla\!\times\!\vec{E})\!+\!\frac{\partial\vec{B}}{\partial t}\!+\!2l_{P}^{2}\theta_{3}\nabla^{2}(\nabla\!\times\!\vec{E})\!-\!2\theta_{8}l_{P}\nabla^{2}\vec{E}=0,$ (6)
@@
The Eq. (II) is an infinite expression. In the literature, it is common to truncate it at the first order in the non-linear term in the magnetic field. When the field equations are calculated, this non-linear term contributes to Eq. (II). The main quantities of the electromagnetic theory have already been calculated and analyzed previously, as it can be seen in Alfaro:2001rb ; Li:2020uef . In the present work, such quantities are not of primary interest; the modified Maxwell equations in the context of LQG as well as its corresponding Hamiltonian are the key equations for the development of the present contribution.
@@ III The Noether Procedure as our path towards the non-Abelian scenario
To formulate a non-Abelian theory for massless spin-1 particles, one may follow different approaches. In this paper, we pick out the so-called Noether method Deser:1969wk ; VanNieuwenhuizen:1981ae , which essentially consists in generating self-interactions among vector fields through the variation of the original free Lagrangian and proposing self-interactions by coupling these vector fields with the Noether current of their initially global symmetry. To realize this program, we start off from an N-plet of vector fields (at the end of the procedure, they will become our Yang-Mills gauge bosons) in an arbitrary representation of the symmetry group as given below:
$A_{m},m=1,2,...,N\ \ \ \ \ \ \ \ \phi_{n},n=1,2,...,N$ (8)
in an N-dimensional representation of the symmetry group. We can define the electric- and magnetic-like fields in terms of the family of the potential field above:
$\vec{E}_{m}=-\nabla\cdot\phi_{m}-\frac{\partial\vec{A}_{m}}{\partial t},$ (9)
and
$\vec{B}_{m}=\nabla\times\vec{A}_{m}.$ (10)
It is important to emphasize that the indices that shall be used from now on are associated with the N-dimensional representation of a general gauge group; they do not correspond any longer to the $su(2)$ mentioned in the previous Section. The symmetry transformations of the potential fields, Eqs. (9), and (10) read as given below:
@@
We are working in the general case of an $SU(N)$ gauge group. So, $w_{h}$ stands for the parameters of the $SU(N)$ transformations, whereas the 3-index symbol $f_{hij}$ represents the structure constants of $SU(N)$. To initiate the Noether procedure, we start off with the free Lagrangian of the theory. To get this Lagrangian, we perform a Legendre transformation by taking Eq. (II) and writing
$L_{LQG}$ $=\frac{1}{Q^{2}}\int d^{3}x\Biggl{\{}\left[1+\theta_{7}\left(\frac{l_{P}}{\mathcal{L}}\right)^{2+2\Upsilon}\right]\frac{1}{2}(\vec{E}^{2}-\vec{B}^{2})$
$+\theta_{3}l_{P}^{2}(\vec{B}\nabla^{2}\vec{B}+\vec{E}\nabla^{2}\vec{E})+\theta_{2}l_{P}^{2}\vec{E}\cdot[\nabla\cdot(\nabla\vec{E})]$
$+\theta_{8}l_{P}[\vec{B}\cdot(\nabla\times\vec{B})+\vec{E}\cdot(\nabla\times\vec{E})]$
$+\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(\vec{B}^{2})^{2}\Biggl{\}},$ (17)
@@
$\vec{j}_{h}\!$ $=\!-f_{hij}\{A_{\gamma}[(\vec{E}_{i}\cdot\phi_{j})\!-\!(\vec{B}_{i}\!\times\!\vec{A}_{j})]-\theta_{3}l_{P}^{2}[(\vec{B}_{i}\!\times\!\nabla^{2}\vec{A}_{j})$
$-(\nabla^{2}\vec{B}_{i}\times\vec{A}_{j})-(\nabla^{2}\vec{E}_{i}\cdot\phi_{j})-(\vec{E}_{i}\cdot\nabla^{2}\phi_{j})]$
$-\theta_{8}l_{P}[2(\nabla^{2}\vec{A}_{i}\times\vec{A}_{j})-(\vec{E}_{i}\times\vec{\dot{A}}_{j})-(\nabla\times\vec{E}_{i})\phi_{j}]$
$+4\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(B^{2}\cdot\vec{B}_{i}\times\vec{A}_{j})\},$ (20)
@@
$j_{h}^{0}$ $=-f_{hij}\{A_{\gamma}(\vec{E}_{i}\cdot\vec{A}_{j})+\theta_{3}l_{P}^{2}[(\nabla^{2}\vec{E}_{i}\cdot\vec{A}_{j})$
$+(\vec{E}_{i}\cdot\nabla^{2}\vec{A}_{j})]-2\theta_{8}l_{P}(\nabla\times\vec{E}_{i})\cdot\vec{A}_{j}\}.$ (21)
@@
The process will be the same, but with an additional step. Before identifying the new conserved currents, we must consult the new field equations derived from the Lagrangian of the previous step. This is so because the new identified currents must be on-shell. We highlight this point, because the on-shell currents should be obey the field equations of the system at each step. From the Lagrangian, we can derive new Gauss’ electricity and Ampère-Maxwell equations. This allows us to find the currents to be coupled to the fields to once again insert in the Lagrangian. Each current is coupled to a field and depends on a coupling parameter $l^{\prime}$, which represents the coupling constant for this new term
@@
Eq. (26) is the last Lagrangian expression obtained, and it is no longer necessary to re-calculate other currents. The reason being that the expression of the new Lagrangian no longer depends on derivatives of the fields. Should the same steps be repeated, we would obtain the same expressions for the currents as before. From Eq. (26), we can obtain the field equations, namely, Gauss for the electric field and Ampère-Maxwell for the magnetic field. Notice that the coupling constants $l$ and $l^{\prime}$ are, in principle, arbitrary parameters of the theory. Thus, by invoking the principle of universality, we can choose them in such a way that there is only a gauge coupling constant, exactly as it is the case for Yang-Mills theories Cheng:1984vwu . We set $l=\frac{1}{4}g$ and $l^{\prime}=\frac{1}{2}g$ and from these two equations of motion, it is possible to define the equations for the electric and magnetic fields in terms of the potentials in the Yang-Mills formalism with LQG contributions:
$\vec{E}_{h}\!$ $=\!A_{\gamma}[-\nabla\phi_{h}\!-\!\vec{\dot{A}}_{h}\!+\!gf_{hij}\phi_{i}\vec{A}_{j}]\!-\!\frac{3g}{2}\theta_{3}l_{P}^{2}f_{hij}[\vec{A}_{i}(\nabla^{2}\phi_{j})$
$+(\nabla^{2}\vec{A}_{i})\!\cdot\!\phi_{j}]+\theta_{3}l_{P}^{2}f_{hij}[(\nabla\vec{A}_{i})\!\cdot\!(\nabla\phi_{j})]$
$+\frac{g}{2}\theta_{8}l_{P}f_{hij}(\vec{A}_{i}\times\vec{\dot{A}}_{j}),$ (27)
@@
$\vec{B}_{h}$ $=A_{\gamma}[(\nabla\times\vec{A}_{h})+\frac{g}{2}f_{hij}(\vec{A}_{i}\times\vec{A}_{j})]+2l_{P}^{2}\theta_{3}\nabla^{2}\vec{B}_{h}$
$-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\cdot[\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-[\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})]]$
$\!+\!\frac{g}{2}\theta_{8}l_{P}f_{hij}[(\vec{\dot{A}}_{i})\!\cdot\!\phi_{j}]\!+\!2\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}[2(\vec{B}^{2}\vec{B}_{h})$
$+3gf_{hij}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}].$ (28)
We call the attention to the fact that the equations above, Eqs. (III) and (III), are the last possible ones to be obtained through our method. The procedure stops here. This is so because, if we repeat the procedure again, we will find the same currents that have been obtained in the previous step. This helps us to understand that the contributions of LQG effects in the electromagnetic sector, together with the non-Abelian formulation, add significant nonlinear terms to the general equations for the electric and magnetic fields, respectively. This is still a contribution arising from the equations of motion, which already contain these derivative terms. Here, we have all the necessary ingredients to obtain all the non-Abelian Maxwell equations, if we were working with the electromagnetism of the Standard Model of Particle Physics. As previously mentioned, the first two equations, i.e, Gauss’s law for electricity and the Ampère-Maxwell equation are obtained from Eq. (26) leaving only two equations yet to be derived. These latter equations can be obtained by taking the curl of Eq. (III) and the divergence of Eq. (III) would be sufficient. However, in electrodynamics with LQG effects, this is not possible, because the Faraday-Lenz equation is modified and, by doing this and taking the Abelian limit, we do not re-obtain the Eqs. (5), (6) as it should be the case. According to the current literature Sorokin:2021tge ; Gaete:2014nda , the non-Maxwellian extensions of electrodynamics do not modify the Faraday-Lenz equation, since together with Gauss’ law for magnetism, these equations are obtained as a consequence of the Bianchi Identities written in terms of the gauge-covariant derivatives. The electrodynamics presented in this paper modifies Eq. (6), which allows us to conclude that the Bianchi Identities are modified. It is therefore necessary to follow another path to attain these two equations; one possible way consists in getting them from the Hamilton-Jacobi equations. By performing a Legendre transformation on the Lagrangian Eq. (26) to find its corresponding Hamiltonian, we can derive the equations for the electric field, $E$, and the vector potential, $A$: